# Axion dark matter from inflation-driven quantum phase transition

### 13 Feb 2024 v1

@@ V Conclusions
Our mechanism allows for a much lower axion decay constant than misalignment, and covers a large range of scales for ultralight DM, from sub-eV down to $10^{-24}~{}{\rm eV}$ (for $N=60$ e-folds, or even lower with more e-folds). The viable parameter space can be probed by future axion experiments, including haloscopes and nuclear clocks, and next-generation CMB surveys as well as $21$ cm telescopes, as shown in Fig. 2. Our mechanism is applicable to both the QCD axion and ALPs. In addition, we expect that it can also be applied to other bosonic ultralight DM scenarios. A more detailed investigation of the phenomenology of our mechanism will be carried out in separate work Ismail et al. .
@@ Acknowledgements.
We would like to thank Haipeng An, Ying-Ying Li, Yuichiro Nakai, Maxim Perelstein, Gilad Perez, Maximilian Ruhdorfer, Ziwei Wang, and Taewook Youn for helpful discussions and feedback. We thank Les Houches workshop PhysTeV 2023 for inspiring us to start this work. AI is supported in part by the NSF grant PHY-2014071 and in part by NSERC, funding reference number 557763. SJL and BY are supported by the Samsung Science Technology Foundation under Project Number SSTF-BA2201-06. SJL is also supported by the National Research Foundation of Korea (NRF) funded by the Korea government (MEST).
@@ Appendix B Isocurvature bound
with $\eta_{\rm i}=f_{\eta}\theta_{\rm i}$ the initial field displacement and $\theta_{\rm i}$ the initial misalignment angle. The latest Planck measurements give $\beta_{\rm iso}<0.038$ Aghanim et al. (2020), which puts an upper bound on the inflation scale:

### 15 Mar 2024 v2

@@ V Conclusions
Our mechanism allows for a much lower axion decay constant than misalignment, and covers a large range of scales for ultralight DM, from sub-eV down to $10^{-24}~{}{\rm eV}$ (for $N=60$ e-folds, or even lower with more e-folds). The viable parameter space can be probed by future axion experiments, including haloscopes, nuclear clocks, CASPEr, and next-generation CMB surveys as well as $21$ cm telescopes, as shown in Fig. 3. Our mechanism is applicable to both the QCD axion and ALPs. In addition, we expect that it can also be applied to other bosonic ultralight DM scenarios. A more detailed investigation of the phenomenology of our mechanism will be carried out in separate work Ismail et al. .
@@ Acknowledgements.
We would like to thank Haipeng An, Abhishek Banerjee, Ying-Ying Li, Yuichiro Nakai, Maxim Perelstein, Gilad Perez, Maximilian Ruhdorfer, Konstantin Springmann, Stefan Stelzl, Xin Wang, Ziwei Wang, and Taewook Youn for helpful discussions and feedback. We thank Les Houches workshop PhysTeV 2023 for inspiring us to start this work. AI is supported in part by the NSF grant PHY-2014071 and in part by NSERC, funding reference number 557763. SJL and BY are supported by the Samsung Science Technology Foundation under Project Number SSTF-BA2201-06. SJL is also supported by the National Research Foundation of Korea (NRF) funded by the Korea government (MEST).
@@ Appendix B Isocurvature bound
with $\eta_{\rm i}=f_{\eta}\theta_{\rm i}$ the initial field displacement and $\theta_{\rm i}$ the initial misalignment angle. The latest Planck measurements give $\beta_{\rm iso}<0.038$ Aghanim et al. (2020), which puts a lower bound on the ratio between PQ breaking scale and inflationary Hubble scale: